4+4-7+17+89+765


1
Expert's answer
2022-05-02T16:04:36-0400

"4+4-7+17+89+765 =\\
8-7+17+89+765=\\
1+17+89+765=\\
18+89+765=\\
107+765=\\
872"

Answer:872